2. Show from national income accounting that
a. An increase in taxes(while transfers remain constant) must imply a change in net exports, government purchases, or the saving-investment balance.
b. An increase in disposable personal income must imply an increase in consumption or increase saving.
c. An increase in both consumption and saving must imply an increase in disposable income.
1
Expert's answer
2017-09-12T02:24:06-0400
а. We determine taxes from the identity for the economy with the state and foreign trade
S - I = (G + TR - TA) + NX
TA = G + TR + NX + I - S
If the tax increases, the other side of the equation should change.
Thus, with a constant transition, the change must be а higher G, higher NX, higher I or lower S.

b. Starting from the conditions of equation YD = S + C, when the disposable income is increased, the change could be saved or consumed.

c. Increase in consumption and savings should mean an increase in disposable income, which is consistent with the above equation.